OpenStudy (anonymous):

solve for x: log$2(x+1)+log$2(3x-5)=log$2(5x-3)+2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(x+1)(3x-5) = 4(5x-3)

OpenStudy (anonymous):

okay......can you keep going?

OpenStudy (anonymous):

3x^2 -2x -5 = 20x -12 3x^2 - 22x +7 = 0

OpenStudy (anonymous):

does the last line factor into (x-7)(3x-1)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x = 7 or 1/3 now substitute both these values into ur initial eqn to double check

OpenStudy (anonymous):

okay and on the first line you wrote 4(5x-3).....how did you get the four? 2^2?

OpenStudy (anonymous):

yes 2 can be written as log 4 to base 2

OpenStudy (anonymous):

and then we use log a + log b =log (ab)

OpenStudy (anonymous):

thanks you. thanks were i kept going wrong. The solutions both work. :)

OpenStudy (anonymous):

no prob miss

OpenStudy (anonymous):

can i ask one more thing? if you are solving for x and you take the square root to isolate x, do you include both the positive and negative answers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

like lets say x^2 = k then x = k AND -k

OpenStudy (anonymous):

in some special situations we can neglect the negative answer..bt generally it is essential to consider both

OpenStudy (anonymous):

thanks you!